OpenStudy (anonymous):

the captian of a team of 11 members is 26 year old and the wicket keeper is 3 yrs older if the age of these 2 are excluded the average age of the remaining players is 1 yr less than the average age of the whole team whats the average age of the team

OpenStudy (anonymous):

There are 11 players. We know that one is 26 and one is 26+3=29 Set x as the sum of the 9 remaining players age. We know that (x+26+29)/11 (the average of the whole team) is 1 year less than x/9 ( the average of the remaining team) x/9 +1=(x+26+29)/11 Multiply by 99 11x+99=9x+495 x=198 However we are asked for the average age of the team: (x+26+29)/11=(198+26+29)/11=23 That is your solution

OpenStudy (anonymous):

ok got it thanks